OpenStudy (anonymous):

What is the y-intercept of the function y = 50(0.03)x? Answer x 1 0.03 50

OpenStudy (anonymous):

@zordoloom the same with this one..

OpenStudy (anonymous):

y = 50*(0.03)^x?

OpenStudy (anonymous):

it exponential

OpenStudy (anonymous):

y = 50*1=?

OpenStudy (anonymous):

The answer is 50?

OpenStudy (anonymous):

yup